# Scaling limits of critical FK-decorated random planar maps with $q=4$

### v1

@@ Appendix B Numerical simulations
Figure 10: Numerical simulations of hamburger-cheeseburger trajectories for $p\in\{0.1,0.4,0.5,0.6,0.9\}$. We took $N=10^{6}$ steps and represented the discrepancy and burger count, both rescaled by $\sqrt{N}$. We see that the trajectory becomes thinner and thinner as $p$ increases. Indeed, the rescaled discrepancy collapses when $p\geq 1/2$ by Sheffield’s result [SheffieldScott2016QGAI]. We stress that there is no a priori monotonicity in $p$ in the inventory accumulation model. At the critical value $p=1/2$, which is the focus of this paper, the $\sqrt{N}$ scaling is only off by a logarithmic factor, which is very subtle to detect from simulations.

### v2

@@ Appendix B Numerical simulations
Figure 10: Numerical simulations of hamburger-cheeseburger trajectories for $p\in\{0.1,0.4,0.5,0.6,0.9\}$. We took $N=10^{6}$ steps and represented the discrepancy and burger count, both rescaled by $\sqrt{N}$. We see that the trajectory becomes thinner and thinner as $p$ increases. Indeed, the rescaled discrepancy collapses when $p\geq 1/2$ by Sheffield’s result [SHE16]. We stress that there is no a priori monotonicity in $p$ in the inventory accumulation model. At the critical value $p=1/2$, which is the focus of this paper, the $\sqrt{N}$ scaling is only off by a logarithmic factor, which is very subtle to detect from simulations.